2/(x-1) = 3/(x+3) + 1/(x+3)(x-1)
What are the first few steps to solving this equation for x?

Answer by stanbon(75887)

2/(x-1) = 3/(x+3) + 1/(x+3)(x-1)
What are the first few steps to solving this equation for x?
----
When you have denominators in an equation, get rid of them:
---
Multiply thru by (x+3)(x-1) to get:
2(x+3) = 3(x-1) + 1
----
2x+6 = 3x-2
x = 8
----------------------
Cheers,
Stan H.
==================